# Mechanisms for producing Primordial Black Holes from Inflationary Models Beyond Fine-Tuning

### v1

@@ 3.2 Step-like potential
3.2.1 An alternative single field mechanism
In the pursuit of models that require minimal fine-tuning, the step-like potential emerges as a promising approach in single field inflationary theories [40, 41]. This potential reads as:
$V(\varphi)=V_{0}\left(1+\frac{1}{2}\sum_{i}A_{i}\left(1+\tanh\left[c_{i}\left(\varphi-\varphi_{i}\right)\right]\right)\right)\,,$ (8)
where $V_{0}$ is fixed for the requirement of the amplitude of power spectrum at CMB scales and $A_{i}$, $c_{i}$, and $\varphi_{i}$ are parameters that define the amplitude, steepness, and position of the steps, respectively. These steps introduce localized features in the potential, enabling significant enhancements in the scalar power spectrum.
A step-like feature in the inflaton potential results in a rapid decrease of potential energy and leads to spectral distortions [40, 41, 77]. Such features, where the potential energy abruptly transitions from one constant value to another at one or more points, can significantly enhance the spectrum over a specific range of scales.
@@ 3.2.2 Studying the fine-tuning
The potential described in Eq. (8) can be integrated into the $\alpha$-attractor framework of supergravity[41]. In our study we have the following form that adheres to the observable constraints of inflation:
$V=(1-e^{-d\varphi})^{2}\left(V_{0}+\sum_{i=0}^{N}B_{i}\left(-1+\tanh[c(\varphi-\varphi_{i})]\right)\right)$ (9)
where $\rm{V_{0}}$, $\rm{c}$, $\rm{d}$ and $\rm{\varphi_{i}}$ are parameters. The parameter $\rm{V_{0}}$ ensures that the potential respect the CMB constraints of inflation and the others identify the steepness and the position of the step as previously.
In this study, we focus on the potential given in Eq. (9) to explore a substantial enhancement of the scalar power spectrum. The following parameters are proposed to achieve such an enhancement at $\mathcal{O}(10^{-2})$ orders of magnitude:
@@
For initial condition we assume $\phi_{ic}=6$. The prediction for spectral index, $\rm{n_{s}}$, and tensor to scalar ratio, $\rm{r}$, are given respectively:
@@
and they respect the observable constraints of inflation [21].
Figure 2: The power spectrum derived from the potential described in Eq. (9) for various choices of the parameter c.
The potential (9) includes four steps necessary to achieve the desired enhancement level. The dependence of the power spectrum’s peak height on the parameter $c$ is depicted in Fig. 2. The selection of this potential facilitates the formation of PBHs at $10^{-6}M_{\odot}$. However, different outcomes can be achieved by varying the choice of the parameter. A proper choice of parameters that can lead to significant results is presented in a subsequent section.
As previously described, we evaluate the parameter $\Delta_{p_{i}}^{P_{R}^{peak}}$, defined in Eq. (1), and present our findings in Table 2. We study the parameters $\rm{B_{i}}$, $\rm{c}$ and $\rm{d}$ which are responsible for the step features. We also examine the distance of the steps as $|\phi_{i+2}-\phi_{i+1}|$ where $\rm{i=0,1}$, etc. As we can notice by the Table 2, the parameter which need more fine-tuning in this case is the distances of the steps $|\phi_{i+2}-\phi_{i+1}|$. However, it is noteworthy that this mechanism exhibits a substantial reduction in the amount of fine-tuning compared to the earlier case of an inflection point in the scalar potential, see Tables 1 and 2.
@@
[…]
Table 2: Analysis of fine-tuning for the potential given in Eq. (9).
@@ 4 Multi-field models
Multi-field inflationary models have been extensively studied for their role in generating PBHs. In this section, we discuss the amount of fine-tuning required in multi-field scenarios, focusing on two distinct mechanisms: one with a non-canonical kinetic term [53, 54, 55, 56, 57, 58, 59] and another with a canonical kinetic term accompanied by a waterfall trajectory [60, 61, 62, 63, 64, 65].
@@ 4.1 Two field model with a non-canonical kinetic term
In this subsection we present a two-field model featuring a non-canonical kinetic term. Although numerous mechanisms have been proposed in the literature, we specifically focus on the one detailed in Ref. [54]. The action for this two-field toy model is expressed as:
@@
In this model the inflaton field $\phi$ interacts with a non-canonical scalar field $\chi$ via an exponential dilaton-like coupling [78]. The potential is defined by:
@@
Here $b_{1}$ is a coupling parameter between the fields and $V_{0}$, $\phi_{0}$, and $m_{\chi}$ are the model parameters. We adopt the parameter choices from [54] as follows
@@
For initial conditions we have:
@@
as they lead to acceptable values for the prediction of spectral index, $\rm{n_{s}}$, and tensor to scalar ratio, r [54, 21].
Figure 3: The power spectrum derived from the Eqs. (13) and (12) for various choices of $b_{1}$ presenting in the kinetic term.
In Fig. 3, we depict the power spectrum for various settings of the parameter $b_{1}$. It is evident that the enhancement of the power spectrum varies with the choice of this parameter. However, the impact of other parameters on the power spectrum also requires examination. Therefore, in Table 3, we present the degree of fine-tuning by evaluating Eq. (1) for all parameters, including the potential parameters $m_{\chi}$ and $\phi_{0}$. It should be noted that the parameter $V_{0}$ does not influence the dynamics of the fields, and for this reason, it is not included in this table.
@@
[…]
Table 3: Analysis of fine-tuning for the potential given in Eqs. (13) and (12).
The results presented in Table 3 indicate that fine-tuning in multi-field inflationary models can be significantly reduced compared to single field inflationary models, particularly those with an inflection point. For a more detailed analysis, it is also essential to examine how the initial conditions of the fields may influence these results.
@@ 4.2 Hybrid model
As an additional mechanism, we study a hybrid model characterized by a waterfall trajectory. It is important to outline some fundamental aspects of this framework. The hybrid model is derived from a globally supersymmetric (SUSY) renormalizable superpotential [79, 80]:
@@
A variety of hybrid models have been explored within the literature to elucidate the production of PBHs [60, 61, 62, 63, 64, 65]. We study the fine-tuning aspects of hybrid models, with a particular focus on the approach outlined in Ref. [62]. We selected this model due to the unique characteristics of its two-field potential, which not only serves as an attractor for the initial conditions of both fields but also ensures that, regardless of these starting conditions, the trajectory consistently leads to the potential’s plateau. Such convergence is crucial as it initiates a slow-roll phase, an important element in the dynamics of the model. The potential is given as follows:
@@
where $M$, $m$, $g$, $d$ and $\psi_{0}$ are parameters. For the parameters we have the following
@@
Hence this model align with the observable constraints of inflation [21, 62].
Figure 4: The power spectrum derived from the Eq. (18) for various choices of parameter $\rm{10^{6}|d|}$
In Fig. 4, we depict the scalar power spectrum derived from the potential in Eq. (18) for various settings of the parameter $d$. This plot illustrates that extensive fine-tuning is not necessary to achieve a significant peak in the power spectrum. As in previous mechanisms of PBHs, we analyze the impact of other parameters on the power spectrum. The parameters $m$ and $g$ influence the position of the power spectrum peak. We assess the amount of fine-tuning using Eq. (1) and present our results in Table 4.
@@
Table 4: Analysis of fine-tuning for the potential given in Eq. (18) .
As one can notice from Table 4, this mechanism significantly reduces the amount of fine-tuning required, compared to both the single field inflationary mechanism and the two-field mechanism with a non-canonical kinetic term. An important notice here is that this mechanism is not dependent on the initial condition of the fields and hence it can provide us with a more natural explanation of PBH production.
@@ 5 Spectator field
From our previous analysis, it is evident that multi-field inflationary models can significantly reduce the fine-tuning required for the production of PBHs. Additionally, the implementation of a spectator field as a solution to this issue has been extensively explored in the literature [67, 68, 69, 70], offering a promising way to circumvent the problem.
We specifically examine the model presented in Ref.[68], where the spectator field experiences quantum stochastic fluctuations. These fluctuations lead the field to assume distinct mean values across different Hubble patches. Considering the vast number of these patches, it is inevitable that some will exhibit conditions allowing the spectator field to attain values necessary for quantum fluctuations to generate significant curvature fluctuations. The potential of the spectator field is described by:
@@
In this work, we further evaluate the amount of fine-tuning previously assessed to models with a spectator field. We employ the fine-tuning measure defined in Eq. (1) as we did it before. This particular mechanism does not require an enhancement in the scalar power spectrum, thereby eliminating the need to consider peaks in our evaluation. Instead, we focus on the fractional abundance of PBHs ($f_{\rm PBH}$) as our primary quantity $Q$. We specifically analyze the parameter $M$, as the parameter $\Lambda$ merely sets the scale of the spectator field and does not influence the dynamics. We found that $\Delta_{M}^{f_{\rm{PBH}}}\sim\mathcal{O}(1)$. Therefore, in this mechanism we can avoid the need of fine-tuning in order to explain the production of PBHs.
We need to remark here that the stochastic nature of the spectator field during inflation plays a crucial role in suppressing fine-tuning. In this mechanism, variations in the model’s single parameter $\rm M$, which is on the same order as the Hubble parameter at CMB scales, can be effectively compensated by the specific inflationary model we study each time. A detailed analysis can be found in [68].
@@ 6 Step-features potential over inflection-point: An additional advantage in implementation of GWs
In this section, we explore the implementation of a step-like feature potential for the production of PBHs, in comparison to models utilizing an inflection point in the potential. We demonstrate how this mechanism can not only lead to a significant enhancement and thus to copious production of PBHs, but also potentially explain the signal of GWs at low frequencies.
Figure 5: The power spectrum derived from the potential given in Eq. (9) with the choice of parameters given in (22).
To begin with, this approach facilitates a significant enhancement of the power spectrum while substantially reducing the need for fine-tuning. The reinforcement of the power spectrum is depicted in Fig. 5. For this plot we assume the following set of parameters:
@@
and for initial condition we have $\phi_{ic}=6$. Hence the prediction of spectral index and tensor-to-scalar ratio is given as follows:
@@
which are in complete accordance with the observable constraints of inflation [21]. We note here that the number of efolds we obtain for this set of parameters is $N_{tot}-N_{CMB}\approx 50$, where $N_{CMB}$ denotes the number of efold at the pivot scale.
For this enhancement the mass of PBHs is about $10^{-1}M_{\odot}$ with the formalism of this evaluation presented in Appendix A.2. It is noteworthy to mention here that recently a few intriguing PBHs in low mass ratio binaries have been reported in GW observations [81, 82]. Therefore, the predicted range of masses of PBHs from this step-like mechanism can be related with the aforementioned searches and then it needs further investigation.
Figure 6: The energy density of GWs for the potential given in Eq. (9) with the choice of parameters given in (22).
In Fig. 6, we depict the energy density of GWs, denoted as $\Omega_{GW}$, using the specific parameters discussed in this section. The calculation of $\Omega_{GW}$ is detailed in Appendix A.3. The gray dots on the plot represent the detection sensitivity of the NANOGrav collaboration. Therefore, this mechanism has the interesting characteristic that it can be imprinted in the GW signals detected by NANOGrav [6, 7, 8].
We observe that, in addition to this enhancement, there are also additional oscillations which could be imprinted in scalar GWs, potentially leading to a characteristic signal [48, 51]. Moreover, we need to mention here that in this work we neglect possible effects of non-Gaussianities [83, 84]. Consequently, this mechanism needs further investigation and should be the focus of a future study.

### v2

@@ 3.2 Step-Like Potential
3.2.1 An Alternative Single-Field Mechanism
In the pursuit of models that require minimal fine-tuning, the step-like potential emerges as a promising approach in single-field inflationary theories Kefala et al. (2021); Dalianis et al. (2021). This potential reads as:
$V(\phi)=V_{0}\left(1+\frac{1}{2}\sum_{i}A_{i}\left(1+\tanh\left[c_{i}\left(\phi-\varphi_{i}\right)\right]\right)+B\phi\right)\,,$ (8)
where $V_{0}$ is fixed for the requirement of the amplitude of power spectrum at CMB scales; $\phi$ is the inflaton field; and $A_{i}$, $B$, $c_{i}$, and $\varphi_{i}$ are parameters that define the amplitude, steepness, and position of the steps, respectively. These steps introduce localized features in the potential, enabling significant enhancements in the scalar power spectrum.
A step-like feature in the inflaton potential results in a rapid decrease in potential energy and leads to spectral distortions Kefala et al. (2021); Dalianis et al. (2021); Dalianis (2023). Such features, where the potential energy abruptly transitions from one constant value to another at one or more points, can significantly enhance the spectrum over a specific range of scales.
The distinctive behavior of step-like features in the potential contrasts markedly with other inflationary mechanisms, such as the inflection point mentioned previously. While the inflection point leads to a temporary stop in the field’s motion, leading to a spike in the power spectrum due to prolonged inflation at specific field values, a step-like feature triggers accelerations and decelerations in the inflaton’s velocity. For more details on the differences between these mechanisms, see Appendix A. Unlike the inflection point, the enhancement in the power spectrum in a step-like potential does not rely on the minimum value of the field’s velocity. For this reason, we anticipate that this mechanism can reduce the amount of fine-tuning required.
@@ 3.2.2 Studying the Fine-Tuning
The potential described in Equation (8) can be integrated into the $\alpha$-attractor framework of supergravity, providing a reliable model that adheres to the observable constraints of the CMB, and can lead to an enhancement in the power spectrum Dalianis et al. (2021). However, this model does not achieve a significant enough enhancement of the power spectrum for the production of PBHs. For this reason, in our study focusing on assessing fine-tuning, we modified the well-known Starobinsky potential to incorporate step-like features. Hence, in our study, we have adopted the following form of the potential that aligns with the observable constraints:
$V(\phi)=(1-e^{-d\phi})^{2}\left(V_{0}+\sum_{i=0}^{N}B_{i}\left(-1+\tanh[c(\phi-\varphi_{i})]\right)\right)$ (9)
where $\rm{V_{0}}$, $\rm{c}$, $\rm{d}$, and ${\varphi_{i}}$ are parameters. The parameters $\rm{V_{0}}$ and $\rm{d}$ ensure that the potential with respect to the CMB constraints of inflation and the others identifies the steepness and the position of the step as previously.
In this study, we focus on the potential given in Equation (9) to explore a substantial enhancement of the scalar power spectrum. The following parameters are proposed to achieve such an enhancement at $\mathcal{O}(10^{-2})$ orders of magnitude:
@@
For the initial condition, we assume $\phi_{ic}=6$. The prediction for the spectral index, $\rm{n_{s}}$, and the tensor to scalar ratio, $\rm{r}$, are given, respectively, as:
@@
and they respect the observable constraints of inflation Akrami et al. (2020).
Potential (9) includes the four steps necessary to achieve the desired enhancement level. The dependence of the power spectrum’s peak height on the parameter $c$ is depicted in Figure 2. The selection of this potential facilitates the formation of PBHs at $10^{-6}M_{\odot}$. In Appendix B, there is a detailed analysis for this mass value and the corresponding fractional abundances of PBH. Different outcomes can be achieved by varying the choice of the parameter. A proper choice of parameters that can lead to significant results is presented in a subsequent section.
Figure 2: The power spectrum derived from the potential described in Equation (9) for various choices of the parameter c.
As previously described, we evaluate the parameter $\Delta_{p_{i}}^{P_{R}^{peak}}$, defined in Equation (1), and present our findings in Table 2. We study the parameters ${B_{i}}$, ${c}$, and ${d}$, which are responsible for the step features. We also examine the distance of the steps as $|\phi_{i+2}-\phi_{i+1}|$, where $\rm{i=0,1}$, etc. As we can notice by Table 2, the parameter that needs more fine-tuning in this case is the distances of the steps $|\varphi_{i+2}-\varphi_{i+1}|$. However, it is noteworthy that this mechanism exhibits a substantial reduction in the amount of fine-tuning compared to the earlier case of an inflection point in the scalar potential, see Tables 1 and 2.
Table 2: Analysis of fine-tuning for the potential given in Equation (9).
@@
$B_{i}/V_{0}$ 18.2
$c$ 3.10
$d$ 53.5
$|\varphi_{i+2}-\varphi_{i+1}|$ $80.6$
As we observe in this section, the distinct dynamics of single-field mechanisms—such as the inflection point and step-like features—lead to localized enhancements in the power spectrum. Models with step-like features in their potentials are characterized by sharper peaks and additional oscillatory patterns, unlike the smoother power spectra seen in other potential forms, such as the previously discussed inflection point. Such localized enhancements are crucial for producing PBHs within narrower mass ranges, thus influencing their distinct characteristics (see Appendix B)
Finally, it is crucial to acknowledge that both of the single-field mechanisms discussed have faced scrutiny regarding one-loop corrections to the evaluation of the power spectrum, as highlighted in studies such as Refs. Kristiano and Yokoyama (2022); Fumagalli et al. (2023); Choudhury et al. (2023a, b); Choudhury (2024); Firouzjahi (2023); Inomata et al. (2023). Consequently, exploring multi-field inflationary models could not only reduce the amount of fine-tuning but also offer a viable alternative for the study of PBHs.
@@ 4 Multi-Field Models
Multi-field inflationary models have been extensively studied for their role in generating PBHs. In the context of multi-field mechanisms, it is anticipated that the degree of fine-tuning required will be lessened. This is expected because the constraints imposed by the CMB power spectrum and the demands for PBH production may not be fully addressed by a single-field. However, within multi-field inflation, various mechanisms may lead to differing degrees of fine-tuning, necessitating further investigation to determine their efficacy. In this section, we discuss the amount of fine-tuning required in multi-field scenarios, focusing on two distinct mechanisms: one with a non-canonical kinetic term Aldabergenov et al. (2022); Braglia et al. (2020); Pi and Sasaki (2023); Gundhi et al. (2021); Geller et al. (2022); Chen et al. (2023); Wang et al. (2024) and another with a canonical kinetic term accompanied by a waterfall trajectory Afzal and Ghoshal (2024); Spanos and Stamou (2021); Braglia et al. (2023); Clesse and García-Bellido (2015); Tada and Yamada (2023); Dimopoulos (2022).
@@ 4.1 Two Field Model with a Non-Canonical Kinetic Term
In this subsection, we introduce a two-field model incorporating a non-canonical kinetic term. The enhancement of the power spectrum in this model arises through a large non-canonical kinetic coupling within a two-field inflationary framework. This mechanism is characterized by a temporary tachyonic instability of the isocurvature perturbations during the transition between two distinct inflationary phases Aldabergenov et al. (2022); Braglia et al. (2020); Gundhi et al. (2021); Geller et al. (2022). Such instability leads to a significant amplification of the curvature perturbations, which can be imprinted in both the production of PBHs and GWs .
Although numerous models with a non-canonical kinetic term have been proposed in the literature, we specifically focus on the one detailed in Ref. Braglia et al. (2020). The action for this two-field toy model is expressed as:
@@
In this model, the inflaton field $\phi$ interacts with a non-canonical scalar field $\chi$ via an exponential dilaton-like coupling Kallosh and Linde (2022). The potential is defined by:
@@
Here, $b_{1}$ is a coupling parameter between the fields and $V_{0}$, $\phi_{0}$, and $m_{\chi}$, which are the model parameters. We adopt the parameter choices from Braglia et al. (2020) as follows:
@@
For the initial conditions, we have:
@@
as they lead to acceptable values for the prediction of the spectral index, $\rm{n_{s}}$, and the tensor to scalar ratio, r Braglia et al. (2020); Akrami et al. (2020).
In Figure 3, we depict the power spectrum for various settings of the parameter $b_{1}$. It is evident that the enhancement of the power spectrum varies with the choice of this parameter. However, the impact of other parameters on the power spectrum also requires examination. Therefore, in Table 3, we present the degree of fine-tuning by evaluating Equation (1) for all parameters, including the potential parameters $m_{\chi}$ and $\phi_{0}$. It should be noted that the parameter $V_{0}$ does not influence the dynamics of the fields, and for this reason, it is not included in this table.
Figure 3: The power spectrum derived from Equations (13) and (12) for various choices of $b_{1}$ present in the kinetic term.
Table 3: Analysis of fine-tuning for the potential given in Equations (13) and (12).
@@
$m_{\chi}$ 1.72
$b_{1}$ 71.7
$\phi_{0}$ 3.02
The results presented in Table 3 indicate that fine-tuning in multi-field inflationary models can be significantly reduced compared to single-field inflationary models, particularly those with an inflection point. For a more detailed analysis, it is also essential to examine how the initial conditions of the fields may influence these results.
@@ 4.2 Hybrid Model
In this section, we delve into a hybrid model characterized by a waterfall trajectory. The key aspect of this mechanism is the tachyonic instability that occurs during the transition between two stages of inflation. This instability results in a significant enhancement of the power spectrum at scales corresponding to the transition. Unlike the previous mechanism where the instability arises from strong kinetic coupling, here the instability is triggered when the inflaton field reaches a critical point, leading the so-called waterfall field to develop a tachyonic solution. In contrast to the previous mechanism requiring intricate parameter setups, this hybrid model naturally derives its instability solely from the field’s potential, enhancing the power spectrum in a more predictable manner. This approach to PBH production was first proposed in Clesse and García-Bellido (2015).
Before presenting a specific model for this mechanism, it is important to outline some fundamental aspects of this framework. The hybrid model is derived from a globally supersymmetric (SUSY) renormalizable superpotential Copeland et al. (1994); Dvali et al. (1994):
@@
A variety of hybrid models have been explored within the literature to elucidate the production of PBHs Afzal and Ghoshal (2024); Spanos and Stamou (2021); Braglia et al. (2023); Clesse and García-Bellido (2015); Tada and Yamada (2023); Dimopoulos (2022). We study the fine-tuning aspects of hybrid models, with a particular focus on the approach outlined in Ref. Braglia et al. (2023). We selected this model due to the unique characteristics of its two-field potential, which not only serves as an attractor for the initial conditions of both fields but also ensures that, regardless of these starting conditions, the trajectory consistently leads to the potential’s plateau. Such convergence is crucial as it initiates a slow-roll phase, an important element in the dynamics of the model. The potential is given as follows:
@@
where $M$, $m$, $g$, $d$, and $\psi_{0}$ are parameters. For the parameters we have the following:
@@
Hence, this model aligns with the observable constraints of inflation Akrami et al. (2020); Braglia et al. (2023).
In Figure 4, we depict the scalar power spectrum derived from the potential in Equation (18) for various settings of the parameter $d$. This plot illustrates that extensive fine-tuning is not necessary to achieve a significant peak in the power spectrum. As in previous mechanisms of PBHs, we analyze the impact of other parameters on the power spectrum. The parameters $m$ and $g$ influence the position of the power spectrum peak. We assess the amount of fine-tuning using Equation (1) and present our results in Table 4.
Figure 4: The power spectrum derived from Equation (18) for various choices of parameter $\rm{10^{6}|d|}$
Table 4: Analysis of fine-tuning for the potential given in Equation (18) .
@@
As one can notice from Table 4, this mechanism significantly reduces the amount of fine-tuning required, compared to both the single-field inflationary mechanism and the two-field mechanism with a non-canonical kinetic term. An important notice here is that this mechanism is not dependent on the initial condition of the fields; hence, it can provide us with a more natural explanation of PBH production.
@@ 5 Spectator Field
From our previous analysis, it is evident that multi-field inflationary models can significantly reduce the fine-tuning required for the production of PBHs. Additionally, the implementation of a spectator field as a solution to this issue has been extensively explored in the literature Kohri et al. (2013); Stamou and Clesse (2024, 2023); Stamou (2024), offering a promising way to circumvent the problem.
We specifically examine the model presented in Ref. Stamou and Clesse (2024), where the spectator field experiences quantum stochastic fluctuations. These fluctuations lead the field to assume distinct mean values across different Hubble patches. Considering the vast number of these patches, it is inevitable that some will exhibit conditions allowing the spectator field to attain values necessary for quantum fluctuations to generate significant curvature fluctuations. The potential of the spectator field is described by:
@@
In this work, we further evaluate the amount of fine-tuning previously assessed for models with a spectator field. We employ the fine-tuning measure defined in Equation (1) as we did before. This particular mechanism does not require an enhancement in the scalar power spectrum, thereby eliminating the need to consider peaks in our evaluation. Instead, we focus on the fractional abundance of PBHs ($f_{\rm PBH}$) as our primary quantity $Q$. We specifically analyze the parameter $M$ as the parameter $\Lambda$ merely sets the scale of the spectator field and does not influence the dynamics. We found that $\Delta_{M}^{f_{\rm{PBH}}}\sim\mathcal{O}(1)$. Therefore, in this mechanism we can avoid the need for fine-tuning in order to explain the production of PBHs.
We need to remark here that the stochastic nature of the spectator field during inflation plays a crucial role in suppressing fine-tuning. In this mechanism, variations in the model’s single parameter $\rm M$, which is of the same order as the Hubble parameter at CMB scales, can be effectively compensated for by the specific inflationary model we study each time. A detailed analysis can be found in Stamou and Clesse (2024).
@@ 6 Step-Features Potential over Inflection-Point: an Additional Advantage in Implementation of Gws
In this section, we explore the implementation of a step-like feature potential for the production of PBHs. We demonstrate how this mechanism can not only lead to a significant enhancement and thus to the copious production of PBHs but also potentially explain the signal of GWs at low frequencies.
To begin with, this approach facilitates a significant enhancement of the power spectrum while substantially reducing the need for fine-tuning. In this section, we consider the potential given in Equation (9) . The reinforcement of the power spectrum is depicted in Figure 5. For this plot, we assume the following set of parameters:
@@
and for initial condition we have $\phi_{ic}=6$. Hence, the prediction of spectral index and tensor-to-scalar ratio is given as follows:
@@
which are in complete accordance with the observable constraints of inflation Akrami et al. (2020). We note here that the number of efolds we obtain for this set of parameters is $N_{tot}-N_{CMB}\approx 50$, where $N_{CMB}$ denotes the number of efolds at the pivot scale.
Figure 5: The power spectrum derived from the potential given in Equation (9), with the choice of parameters given in (22).
For this enhancement, the mass of PBHs is about $10^{-1}M_{\odot}$, with the formalism of this evaluation presented in Appendix B. It is noteworthy to mention here that recently a few intriguing PBHs in low mass ratio binaries have been reported in GW observations Abbott et al. (2023); Phukon et al. (2021). Therefore, the predicted range of masses of PBHs from this step-like mechanism can be related to the aforementioned searches, but then it needs further investigation.
In Figure 6, we depict the energy density of GWs, denoted as $\Omega_{GW}$, using the specific parameters discussed in this section. The calculation of $\Omega_{GW}$ is detailed in Appendix C. The gray dots on the plot represent the detection sensitivity of the NANOGrav collaboration. Therefore, this mechanism has the interesting characteristic that it can be imprinted in the GW signals detected by NANOGrav Agazie et al. (2023a, b); Afzal et al. (2023). It is important to note that the parameter values examined in Figure 6 do not cover the energy densities of GWs observed at larger frequencies, such as those from LVK collaboration Abbott et al. (2016a, b, 2017a, 2017b, 2017c); Dwyer et al. (2015) or those expected from future space-based experiments like LISA Amaro-Seoane et al. (2017) and DECIGO Seto et al. (2001) or ground-based observatories such as ET Maggiore et al. (2020). However, adjusting the parameter scans may yield energy densities that can explain these larger-scale signals. For a detailed review about the induced GWs and the frequency regions of future and present experiments, see Ref. Domènech (2021).
Figure 6: The energy density of GWs for the potential given in Equation (9), with the choice of parameters given in (22).
Finally, we observe that, in addition to this enhancement, depicted in Figure 5, there are also additional oscillations that could be imprinted in the energy densities of the induced GWs shown in Figure 6. These oscillations can potentially lead to a characteristic signal Fumagalli et al. (2021); Mavromatos et al. (2022). Moreover, we need to mention here that in this work we neglect the possible effects of non-Gaussianities Cai et al. (2019); Perna et al. (2024); Liu et al. (2024); Namjoo et al. (2013). Consequently, this mechanism needs further investigation and should be the focus of a future study.